NEW HIGH RESOLUTION ANALYSIS OF H_{2}CO IN THE 3.6 and 4.3 μm REGION BY FOURIER TRANS-FORM SPECTROSCOPY

\underline{A. PERRIN}, J.-M. FLAUD, Laboratoire de Photophysique Moléculaire, CNRS, Université Paris Sud, Campus d’Orsay, Bat 210, 91405 Orsay Cedex, France; A. VALENTIN, Laboratoire de Physique Moléculaire et Applications, C.N.R.S., Université Pierre et Marie Curie, Tour 13, 4 Place Jussieu, 75252 Paris, Cedex 05, France; L. R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, California, USA.

Using new Fourier transform spectra recorded at high resolution at LPMA Paris, it has been possible to perform a new study of the formaldehyde absorption spectra in the 2600-3000cm^{−1} and in the 2200-2500cm^{−1} spectral ranges. This analysis was started using the results obtained by previous studies performed in the same spectral regions ^{a} ^{b}. In the high frequency range, the analysis of the strongest bands, namely ν_{1} and ν_{5} (symmetric and antisymmetric stretching modes) was complicated by the existence of Fermi-type resonances and by A-type, B-type and C-type Coriolis interactions involving the ν_{1} and ν_{5} energy levels and levels from various overtone or combination states namely ν_{2} + ν_{6}, 2ν_{3}, ν_{2} + ν_{4}, ν_{3} + ν_{6} and ν_{3} + ν_{4}. In the low frequency range, which involve the weak 2ν_{4} and 2ν_{6} bands and the very weak ν_{4} + ν_{6} band, the 2ν_{4} ↔ 2ν_{6} Fermi type resonance and the 2ν_{4} ↔ ν_{4} + ν_{6} and 2ν_{6} ↔ ν_{4} + ν_{6} A-type Coriolis interaction had to be considered. For each analysed spectral region, a preliminary calculation of the energy levels was performed taking into account the observed resonances. It appears that these two spectral domains should \underline{not} be considered as independent because of the existence of additional resonances linking levels measured in each spectral region.

aL. R. Brown, R. H. Hunt, and A. Pine, J. Mol. Spectrosc. \underline{75}, 406 (1979).

bF. Ito, T. Nakanaga, and H. Takeo, Spectrochimica Acta. \underline{50A}, 1397 (1994).